the base of a rectangle is 7 inches more than its height. if the area of the rectangle is 30 square inches, find its base and its height
Answer by richwmiller(17219)

h+7=b
2b+2h=30
2h+2*(h+7)=30
4h+14=30
4h=16
h=4
b=11